OpenStudy (anonymous):

Thomas wants to save money for a vacation. Thomas invests $400 in an account that pays an interest rate of 4%. How many years will it take for the account to reach $3,400? Round your answer to the nearest hundredth.

OpenStudy (anonymous):

use the equation i gave you...

OpenStudy (anonymous):

which i have no idea what to plug in where

OpenStudy (anonymous):

Do you know what you are compounding by? Monthly, yearly, continuosly?

OpenStudy (anonymous):

yearly i believe

OpenStudy (anonymous):

if that is the case then A=P(1-R/n)^rt therefore 3400 = 400(1- .04/1)^(.04)t Solve for t